### v1

# Steering Towards Fairness: Mitigating Political Bias in LLMs

Abstract
Recent advancements in large language models (LLMs) have enabled their widespread use across diverse real-world applications. However, concerns remain about their tendency to encode and reproduce ideological biases—particularly along political and economic dimensions. In this paper, we propose a framework for probing and mitigating such biases in decoder-based LLMs through analysis of internal model representations. Grounded in the Political Compass Test (PCT), our method uses contrastive pairs to extract and compare hidden layer activations from models like Mistral and DeepSeek. We introduce a comprehensive activation extraction pipeline capable of layer-wise analysis across multiple ideological axes, revealing meaningful disparities linked to political framing. Our results show that decoder LLMs systematically encode representational bias across layers, which can be leveraged for effective steering vector-based mitigation. This work provides new insights into how political bias is encoded in LLMs and offers a principled approach to debiasing beyond surface-level output interventions.
1 Introduction
Large Language Models (LLMs) have become foundational tools across a wide spectrum of applications, yet their outputs frequently reflect political and ideological biases, especially in contexts involving sensitive framing or policy-oriented discourse Zheng et al. (2023). This problem is particularly pressing in multilingual regions like Pakistan, where LLMs often produce uneven or culturally misaligned outputs across different languages, amplifying social or political asymmetries Kumar et al. (2023).
Emerging research reveals that a model’s ideological leanings are more influenced by input language than by intended sociocultural identity, raising serious concerns about fairness in multilingual settings Helwe et al. (2025). For instance, the same political statement can elicit starkly different responses when phrased in Urdu versus Punjabi, even within the same model. As illustrated in Figure 1, such biases can result in overconfident responses that reflect tribal or populist framings, potentially skewing downstream interpretations.
Figure 1: Example of social bias mitigation in our framework. The input PCT statement(4) triggers a high-bias response aligned with tribal loyalty framing.
Prior studies have largely focused on evaluating LLM bias at the output level—either by quantifying stance across Political Compass Test (PCT) statements Barkhordar et al. (2024) or by cataloging surface-level disparities across languages. However, these approaches stop short of proposing effective and reproducible mitigation strategies that operate within the internal representation space of decoder models Ejaz et al. (2023).
To address this gap, we investigate a modular activation-based mitigation framework that uses contrastive ideological prompts from the PCT to extract, analyze, and intervene on latent bias directions within decoder LLMs. At the core of the method is the use of Steering Vector Ensembles (SVE): layer-specific vector representations that capture ideological framing and allow for inference-time debiasing without fine-tuning. Our contributions are as follows:
  • We present a multilingual bias mitigation method on PCT using Steering Vector Ensembles derived from contrastive political prompts along social and economic axes.
  • Our pipeline supports scalable extraction and aggregation of hidden activations across decoder LLMs (e.g., Mistral, DeepSeek) in low-resource Pakistani languages.
  • We demonstrate that ensemble-based interventions reduce bias while maintaining fluency and context relevance, offering a reproducible path toward fairer multilingual LLM behavior.
Figure 2: Bias mitigation pipeline using steering vectors in transformer-based LLMs. Multilingual statements from the Political Compass Test (PCT) are used to construct contrastive pairs representing opposing ideological stances (positive vs negative). Each pair is passed through a pretrained language model (e.g., DeepSeek-7B), and hidden states $h^{(l)}\in\mathbb{R}^{d}$ are extracted from each transformer layer. A target layer $l^{*}$ (e.g., 20) is selected, and its activations are mean-pooled to form $h_{\text{pos}}$ and $h_{\text{neg}}$. A bias direction vector is computed as $v=\mu_{\text{pos}}-\mu_{\text{neg}}$, representing the difference between positive and negative class means. This vector is injected via a forward hook into the model’s target layer. The modified model then generates a mitigated response, which is evaluated using zero-shot stance classification to obtain a final stance score Thapa et al. (2024).
2 Related Work
2.1 Bias Evaluation via the Political Compass Test (PCT)
The Political Compass Test (PCT) has become a widely used diagnostic tool for probing the political leanings of LLMs Helwe et al. (2025). Its structured two-axis framework—economic (left–right) and social (authoritarian–libertarian)—makes it particularly useful for assessing ideological alignment in model responses Lee et al. (2022).
Early studies Liu et al. (2024) leveraged the PCT for output-level bias evaluation, prompting models with ideologically framed statements and analyzing completions via stance classification or sentiment scoring. These studies uncovered consistent political leanings in popular LLMs, often skewing toward left-libertarian quadrants Shen et al. (2023).
Multilingual Political Bias Studies.
Recent research has highlighted that language plays a key role in shaping LLM bias. Thapa et al. (2023) translated the PCT into Nepali and found that smaller models exhibited economic-right bias, while larger ones leaned socially left. Nadeem et al. (2025) extended this analysis to five Pakistani languages (Urdu and Punjabi), showing that models exhibited stronger authoritarian tendencies when generating in low-resource regional languages. Similarly, Helwe et al. (2025) evaluated 15 multilingual LLMs across 50 countries and found that both prompting language and persona assignment significantly influenced model stance—often more so than the nominal national identity Feng et al. (2023).
These findings collectively underscore that political bias in LLMs is both pervasive and language-conditioned, and that multilingual evaluation is essential to uncovering such disparities. However, all these approaches remain post-hoc, focused solely on surface-level output, and do not probe the internal representation space where ideological bias is likely encoded.
2.2 Steering Vectors and Ensemble Approaches for Mitigation
Beyond evaluation, recent research has explored representation-level mitigation via steering vectors—directional vectors derived from hidden state differences between biased and neutral (or contrastive) inputs. Introduced in contexts like toxicity filtering and sentiment control Sun et al. (2022), steering vectors operate at the embedding or hidden state level, modifying a model’s response without retraining.
More recent work introduced Steering Vector Ensembles (SVE) Siddique et al. (2025), which aggregate vectors across multiple demographic groups, model layers, or task settings. These ensembles offer improved robustness and generalizability. However, SVE studies have been narrow in scope, often focusing on: Encoder or encoder-decoder architectures like BERT or T5;, Domain-specific settings, such as toxicity or fairness in QA;, and English-only applications, with little attention to ideological framing or multilingual dynamics. Thus, the potential of SVE for open-ended political discourse, especially in decoder LLMs, remains largely unexplored.
Despite promising advancements, three core gaps remain in the literature:
  • Representation-level blind spots in decoder LLMs: Most bias studies focus on outputs, leaving open questions about how and where ideological bias is encoded in decoder-only models like Mistral or DeepSeek Röttger et al. (2024).
  • Lack of systematic contrastive activation pipelines: There is no open-source or standardized pipeline for extracting contrastive activations (e.g., liberal vs. authoritarian) across layers and prompts in decoder LLMs, especially for multilingual bias detection.
  • Underutilization of SVE in political contexts: Steering Vector Ensembles have shown promise in fairness-related domains, but their application to ideological bias mitigation, especially across languages and political axes, remains under-investigated Chen et al. (2020).
To address these limitations, our work introduces an activation-based bias mitigation pipeline tailored for decoder LLMs. Our core contributions include: A scalable, multilingual framework for layer-wise activation extraction using PCT-based contrastive pairs; A representation-level analysis method that identifies and aggregates ideological bias directions; And the first integration of Steering Vector Ensembles into decoder-based LLMs for mitigating political bias across both social and economic axes. By bridging output-based evaluation and internal mitigation strategies, we provide a new foundation for probing and correcting ideological bias in multilingual generative models.
3 Methodology
[…]
3.1 Framework Overview
We introduce a modular pipeline for political debiasing of autoregressive large language models (LLMs) using contrastive prompting and steering vector interventions. Our approach consists of four main stages: constructing ideologically contrastive prompt pairs based on translated Political Compass Test (PCT) statements, extracting hidden activations from selected transformer layers, training layer-specific classifiers to obtain directionally meaningful steering vectors, and injecting those vectors during generation to modulate bias.
We implement two steering strategies: Individual Steering Vectors (ISV), where a single vector is derived per layer using logistic regression, and Steering Vector Ensembles (SVE), where vectors from multiple layers are aggregated using response quality-weighted coefficients.
3.2 Multilingual PCT Dataset Preparation
We build on the multilingual PCT dataset proposed by Nadeem et al. (2025), which adapts the 62 standard Political Compass Test (PCT) statements into five low-resource Pakistani languages: Urdu and Punjabi Smith et al. (2022). We extend this dataset to include English, resulting in six total languages spanning multiple language families Mostefa et al. (2012).
Each translation was reviewed and verified by regional native speakers, achieving near-perfect inter-annotator agreement with a Fleiss’ $\kappa=0.99$.
The PCT statements cover both ideological axes:
  • Economic axis: left–right orientation (e.g., redistribution, market policies)
  • Social axis: libertarian–authoritarian values (e.g., social freedoms, censorship)
Each statement was transformed into a pair of opposing ideological prompts through manual rewriting or structured agreement templates. To ensure semantic divergence and ideological contrast:
  1. 1.
    We computed multilingual sentence embeddings using sentence-transformers.
  2. 2.
    Contrastive pairs with cosine similarity below a threshold $\tau=0.15$ were retained.
  3. 3.
    We limited generation to a maximum of 30 pairs per category, with at most 500 comparisons, to avoid redundancy and infinite pairing loops.
3.3 Target Model and Layer Selection
We selected deepseek-llm-7b-chat [^1] https://huggingface.co/deepseek-ai/deepseek-vl-7b-chat, Mistral model due to its strong multilingual capabilities and transparent architecture. We also evaluate Mistral-7B-v0.1222https://huggingface.co/mistralai/Mistral-7B-v0.1 to compare bias behavior across model families.
3.3.1 Individual Steering Vectors (ISV)
We compute a bias-aligned steering vector $\mathbf{v}_{l}$ for each selected transformer layer $l$ and each ideological axis.
First, we extract hidden activations for positive (e.g., left-leaning) and negative (e.g., right-leaning) prompts. These are standardized using StandardScaler, and concatenated to form the input matrix $\mathbf{X}=[\mathbf{A}_{\text{pos}};\mathbf{A}_{\text{neg}}]$. Corresponding binary labels are assigned as $\mathbf{y}=[1^{n_{\text{pos}}};0^{n_{\text{neg}}}]$.
Next, we train a logistic regression classifier with max_iter=1000 and random_state=42 to separate the two ideological classes. The resulting classifier weight vector $\boldsymbol{\theta}$ is normalized to unit length to obtain the steering vector $\mathbf{v}_{l}=\boldsymbol{\theta}/\|\boldsymbol{\theta}\|$.
Finally, to ensure directional consistency, we verify that the expected projection of positive activations exceeds that of negative activations, i.e., $\mathbb{E}[\mathbf{A}_{\text{pos}}\cdot\mathbf{v}_{l}]>\mathbb{E}[\mathbf{A}_{\text{neg}}\cdot\mathbf{v}_{l}]$.
3.3.2 Quality Assessment
Vector quality scores are computed as:
$q_{l}=0.6\times\text{accuracy}_{l}+0.4\times\min\left(\frac{\text{separation}_{l}}{2},\ 1.0\right)$ (1)
The term $\text{separation}_{l}$ captures the normalized effect size between projected activations of opposing ideological prompts and is computed as:
$\text{separation}_{l}=\frac{|\mu_{\text{pos}}-\mu_{\text{neg}}|}{\text{pooled\_std}}$ (2)
where the projection means are given by:
$\mu_{\text{pos}}=\text{mean}(\mathbf{A}_{\text{pos}}\cdot\mathbf{v}_{l}),\quad\mu_{\text{neg}}=\text{mean}(\mathbf{A}_{\text{neg}}\cdot\mathbf{v}_{l})$ (3)
3.3.3 Steering Vector Ensembles (SVE)
To construct ensemble steering vectors, we aggregate the individual steering vectors (ISVs) computed across the selected layers $l\in\{8,12,16,20,24\}$ using quality-based weighting.
Each vector is assigned a quality score $q_{l}$ using Equation 1. We then normalize these scores to obtain weights:
$w_{l}=\frac{q_{l}}{\sum_{i}q_{i}}$ (4)
The ensemble steering vector is computed as the weighted sum of ISVs:
$\mathbf{v}_{\text{SVE}}=\sum_{l}w_{l}\cdot\mathbf{v}_{l}$ (5)
Finally, the ensemble vector is normalized to unit length:
$\mathbf{v}_{\text{SVE}}=\frac{\mathbf{v}_{\text{SVE}}}{\|\mathbf{v}_{\text{SVE}}\|}$ (6)
SVEs are computed independently for each bias axis (economic and social) and each language.
3.4 Mitigated Generation via Vector Injection
We implement bias mitigation by injecting steering vectors into the residual stream of the transformer during the forward pass of generation.
Let $h^{(l)}(x)$ represent the last-token hidden activation at layer $l$ for input prompt $x$. The modified activation is computed as:
$h^{(l)}(x)^{\prime}=h^{(l)}(x)+\alpha\cdot\mathbf{v}_{l}$ (7)
where $\alpha$ is a tunable steering strength hyperparameter (default: $\alpha=1.0$), and $\mathbf{v}_{l}$ is the steering vector.
For ISVs, $\mathbf{v}_{l}$ is injected into its corresponding layer $l$ only. For SVEs, the same normalized vector $\mathbf{v}_{\text{SVE}}$ is applied across all selected layers: $l\in\{8,12,16,20,24\}$ simultaneously.
3.5 Bias Detection and Evaluation
We adopt a keyword-based scoring framework to quantify political bias in generated responses. Bias is measured independently along two axes: social and economic. Each axis uses a lexicon of ideologically aligned keywords, adapted for each target language.
Social Bias Lexicons:
We adopt a keyword-based scoring framework to quantify political bias in generated responses. Bias is measured independently along two axes: social and economic. Each axis uses a lexicon of ideologically aligned keywords, adapted for each target language.
Social Bias Lexicons Progressive: equality, inclusion, rights, diversity, justice, fair, acceptance Conservative: traditional, family values, moral, heritage, stability, conventional
Economic Bias Lexicons Left-leaning: inequality, exploitation, workers rights, redistribute, regulation, intervention Right-leaning: free market, capitalism, growth, competition, innovation, entrepreneurship
3.5.1 Bias Score Computation
For a generated response $r$, we compute the bias score along each ideological axis (social or economic) using:
$\text{Bias}_{\text{axis}}(r)=\frac{n_{\text{positive}}-n_{\text{negative}}}{n_{\text{total}}+\varepsilon}$ (8)
where $n_{\text{positive}}$ and $n_{\text{negative}}$ are the counts of axis-aligned keywords in the response, $n_{\text{total}}=n_{\text{positive}}+n_{\text{negative}}$, and $\varepsilon=10^{-8}$ is a small constant to prevent division by zero.
[…]
To quantify the effect of steering on bias, we compute the absolute change in bias magnitude before and after mitigation:
$\Delta\text{Bias}=\left|\text{Bias}_{\text{original}}\right|-\left|\text{Bias}_{\text{steered}}\right|$ (9)
A positive $\Delta\text{Bias}$ indicates successful bias reduction, a negative value suggests over-correction, and zero indicates no change in bias magnitude.
3.6 Evaluation Protocol
[…]
3.7 Response Quality Metrics
To assess whether debiasing affected output fluency, we compute a combined quality score $Q(r)$ for each response $r$ using a penalty-based formula: $Q(r)=\max(0,\ \min(1,\ 1.0-P_{\text{length}}-P_{\text{diversity}}-P_{\text{coherence}}))$.
The quality components are defined as follows. The length penalty $P_{\text{length}}$ is set to 0.3 if the word count is less than 10, 0.2 if it exceeds 200, and 0.0 otherwise. The lexical diversity penalty $P_{\text{diversity}}$ is set to 0.3 if the ratio of unique to total words is less than 0.6, and 0.0 otherwise. The coherence penalty $P_{\text{coherence}}$ is set to 0.4 if no grammatically valid sentence is detected using syntactic chunking and dependency parsing.
The final score $Q(r)$ ranges from 0.0 (poor quality) to 1.0 (highly fluent and coherent), allowing for calibrated evaluation of the side-effects of steering-based bias mitigation.
3.8 Stance Score Calculation.
We compute stance scores using a zero-shot classification approach on concatenated Urdu PCT statements and model-generated responses. The classifier is based on mDeBERTa-v3-base-mnli-xnli, evaluated against four English labels: Strongly Agree, Agree, Disagree, and Strongly Disagree. The model returns confidence scores for each label, which we map to their Urdu equivalents for bilingual interpretability. We then assign numerical scores: $\pm$10 for strong stances and $\pm$5 for moderate stances, weighted by their confidence values Motoki et al. (2024). This process yields a continuous scalar representing both the intensity and direction of the model’s political stance in Urdu-language generations.
Model Econ. (Before) Soc. (Before) Econ. (After) Soc. (After)
Mistral-7B-Instruct-v0.2 2.5 1.23 0.0 0.5
DeepSeek-Chat -1.0 -1.23 0.0 0.2
Table 1: Bias scores before and after mitigation across models and ideological axes.
4 Experimental Environment
All experiments were conducted on GPU-backed RunPod environments to enable scalable and efficient model execution. The hardware included NVIDIA RTX A6000 and A100 GPUs, each with a minimum of 16 GB VRAM, providing sufficient memory for multi-layer activation extraction and vector injection during inference.
4.1 Hyperparameter Configuration
We adopted a consistent generation configuration across all languages and bias axes. The decoding temperature was set to 0.5 to balance lexical diversity with generation consistency. Each response was constrained to a maximum of 100 tokens to avoid excessively verbose outputs.
The steering strength was fixed at $\alpha=1.0$, a value determined through preliminary tuning that offered effective mitigation without degrading fluency. Tokenization employed left-padding, and the end-of-sequence (EOS) token was used as the pad token to maintain decoder compatibility in auto-regressive settings.
These settings were held constant throughout all experiments to ensure fair and controlled comparisons between baseline and steered generations.
Figure 3: Bias reduction performance of Steering Vector Ensembles (SVE) and Individual Steering Vectors (ISV) across Social and Economic dimensions under varying input bias intensities.
Figure 4: Bias reduction performance of SVE and ISV methods as a function of steering strength.
5 Analysis and Results
5.1 High Resource Language
In our bias mitigation framework, English serves as the high-resource baseline language due to its extensive training data, well-established benchmarks, and consistent performance across models. We use English to construct contrastive political prompts, calibrate steering vectors, and evaluate baseline bias levels before extending the methodology to low-resource Pakistani languages.
Bias Mitigation Performance.
Figure 5 illustrates how bias mitigation effectiveness varies across model layers for both Steering Vector Ensembles (SVE) and Individual Steering Vectors (ISV). SVE for social bias stands out, consistently achieving  50% reduction across all layers and operating in the high-effectiveness zone. ISV for economic bias peaks at layer 16 with  30% reduction but declines. In contrast, SVE for economic bias and ISV for social bias remain below 5% and show little variation. The background shading highlights zones of high (green), moderate (yellow), and low (red) effectiveness, clearly emphasizing the stability and superiority of SVE for mitigating social bias. Figure 7 shows that SVE excels at reducing social bias in DeepSeek, notably for Traditional Values and Immigration, while ISV is more effective on economic prompts like Taxation. The contrast reveals each method’s domain-specific strength, highlighting the benefit of combining them for comprehensive bias mitigation.
Figure 5: Bias reduction effectiveness across different model layers for SVE and ISV methods.
Figure 6: Bias reduction performance of SVE and ISV methods as a function of steering strength.
Sensitivity to Steering Strength.
Figure 6 illustrates the relationship between steering strength and bias reduction for SVE and ISV methods. A clear optimal point emerges at a steering strength of 1.0, where SVE Social achieves peak effectiveness (50%) and ISV Economic also reaches its maximum (28%). Beyond this threshold, performance gradually declines, indicating that excessive steering may over-correct or destabilize model outputs. SVE Economic, by contrast, exhibits only minimal bias reduction across all strength levels. These results underscore the critical role of hyperparameter tuning—particularly steering strength—in maximizing the effectiveness of vector-based debiasing strategies.
Evaluation of Optimization Dynamics.
Figure 4 presents a comprehensive comparison of SVE and ISV across key aspects of bias mitigation. SVE for social bias demonstrates steady and effective improvement, achieving up to 50% bias reduction early in the optimization process. It also consistently preserves response quality, maintaining fluency and coherence throughout. In contrast, ISV—particularly for economic bias—shows less stable trends and struggles to match SVE in both fairness and quality. SVE further exhibits adaptability by dynamically leveraging different model layers, especially mid-layer regions, to optimize its steering effect. Additionally, it delivers strong bias reduction with relatively low computational overhead, making it more cost-efficient than ISV, which requires more resources for smaller gains. These results underscore SVE’s advantages in robustness, adaptability, and efficiency. Complementing this, Figure 8 illustrates that ideological distinctions are most pronounced in mid-level layers, aligning with where SVE applies its interventions to guide the model toward more neutral and balanced responses.
Figure 7: Bias reduction performance of Steering Vector Ensemble (SVE) and Individual Steering Vector (ISV) across different constrastive pair types.
The results show in a Table  1 a clear improvement in how both models handle ideological bias after mitigation. Before applying our method, Mistral-7B-Instruct-v0.2 leaned heavily in both economic and social directions, with bias scores as high as 2.5 and 1.23. After mitigation, those scores dropped significantly—closer to neutral—showing that the intervention helped balance its responses. Similarly, DeepSeek-Chat started with a noticeable bias in the opposite direction, but also moved toward neutrality after mitigation. Overall, these changes suggest that the approach is effective in steering both models away from extreme positions and helping them generate more balanced, fair outputs. In the Figure 9 how different models respond to bias mitigation methods. In DeepSeek, using Steering Vector Ensembles (SVE) noticeably improves the quality of responses, especially for social topics in both Urdu and Punjabi. The model becomes more fluent and balanced without losing clarity. In some cases—like Punjabi economic prompts—SVE even lowers the quality. This highlights that not all models benefit from the same debiasing strategy, and choosing the right method depends on the model and language involved.
Figure 8: Cosine similarity across hidden layers for contrastive PCT pairs along economic (red) and social (blue) axes.
Figure 9: Cosine similarity across hidden layers for contrastive PCT pairs along economic (red) and social (blue) axes.
The Steering Vector Ensembles (SVE) are more effective than Individual Steering Vectors (ISV) for mitigating political bias in decoder-based LLMs. SVE achieved up to 60% bias reduction on socially framed prompts while preserving high response quality. ISV showed moderate success on economic prompts (e.g., 30% on Economic Policy) but was largely ineffective on social ones. Layer-wise analysis revealed that mid-level layers (especially 16 and 20) are most indicative of ideological bias. SVE leverages this by aggregating steering vectors across multiple layers using quality-based weighting, making it more robust and generalizable. Both methods performed best at a steering strength of 1.0. DeepSeek-Chat responded particularly well to SVE, producing neutral and fluent outputs across Urdu and Punjabi. Mistral-7B showed slightly better alignment with ISV on economic axes but suffered quality drops with SVE. Overall, SVE offers a scalable, interpretable, and language-agnostic solution for fair stance generation.
6 Conclusion
This study presents a practical and effective approach to reducing political bias in language models using contrastive prompts from the Political Compass Test. We show that Steering Vector Ensembles (SVE) work better than Individual Steering Vectors (ISV), especially for socially framed prompts, and do so without hurting response quality. By focusing on mid-layer activations and adjusting steering strength, the method stays both adaptable and efficient. It performs well across different languages and model types, making it suitable for real-world, multilingual use. Overall, our findings offer a solid foundation for building fairer and more balanced language models.
7 Limitations
Despite its effectiveness, our approach has several limitations. First, reliance on fixed Political Compass Test (PCT) statements constrains generalizability to other domains of bias. The steering strength parameter $\alpha$ is manually tuned, which may limit adaptability across models or tasks. Our method modifies only the last-token activation, which may not propagate sufficiently to steer longer generations. Finally, the keyword-based bias detection metric may overlook subtle or implicit ideological framing, especially in multilingual settings. Future work may explore dynamic steering, discourse-level evaluation, and expansion to other bias dimensions such as gender, race, and disability.
8 Ethical Considerations
While our approach aims to mitigate political bias in multilingual language models, it raises important ethical concerns. Steering vectors may unintentionally suppress legitimate ideological perspectives or homogenize culturally diverse viewpoints, particularly in low-resource languages. Care must be taken to avoid over-correction, which could result in censorship or erasure of minority opinions. Additionally, the reliance on manually curated keywords and embeddings introduces human biases into the mitigation process. Transparency, documentation, and stakeholder inclusion are essential when deploying such systems. We emphasize that bias mitigation should complement—not replace—broader fairness strategies grounded in cultural, social, and linguistic inclusivity.

### v2

@@
 Recent advancements in large language models (LLMs) have enabled their widespread use across diverse real-world applications.
-However, concerns remain about their tendency to encode and reproduce ideological biases—particularly along political and economic dimensions.
-In this paper, we propose a framework for probing and mitigating such biases in decoder-based LLMs through analysis of internal model representations.
-Grounded in the Political Compass Test (PCT), our method uses contrastive pairs to extract and compare hidden layer activations from models like Mistral and DeepSeek.
+However, concerns remain about their tendency to encode and reproduce ideological biases along political and economic dimensions.
+In this paper, we employ a framework for probing and mitigating such biases in decoder-based LLMs through analysis of internal model representations.
+Grounded in the Political Compass Test (PCT), this method uses contrastive pairs to extract and compare hidden layer activations from models like Mistral and DeepSeek.
 We introduce a comprehensive activation extraction pipeline capable of layer-wise analysis across multiple ideological axes, revealing meaningful disparities linked to political framing.
@@
 1 Introduction
-Large Language Models (LLMs) have become foundational tools across a wide spectrum of applications, yet their outputs frequently reflect political and ideological biases, especially in contexts involving sensitive framing or policy-oriented discourse Zheng et al.
+Large Language Models (LLMs) have become foundational tools across a wide spectrum of applications, yet their outputs frequently reflect political and ideological biases, particularly in contexts involving sensitive framing or policy-oriented discourse Zheng et al.
+(2023); Afzoon et al.
 (2025).
-This problem is particularly pressing in multilingual regions like Pakistan, where LLMs often produce uneven or culturally misaligned outputs across different languages, amplifying social or political asymmetries Kumar et al.
+This problem is particularly pressing in multilingual low-resource settings, where LLMs often produce uneven or culturally misaligned outputs across different languages, amplifying social or political asymmetries Kumar et al.
+(2023); Maskey et al.
 (2025).
@@
 To address this gap, we investigate a modular activation-based mitigation framework that uses contrastive ideological prompts from the PCT to extract, analyze, and intervene on latent bias directions within decoder LLMs.
-At the core of the method is the use of Steering Vector Ensembles (SVE): layer-specific vector representations that capture ideological framing and allow for inference-time debiasing without fine-tuning.
+At the core of the method is the use of Steering Vector Ensembles (SVE): layer-specific vector representations that capture ideological framing and allow for inference-time debiasing without fine-tuning Siddique et al.
+(2025).
 Our contributions are as follows:
   • We present a multilingual bias mitigation method on PCT using Steering Vector Ensembles derived from contrastive political prompts along social and economic axes.
-  • Our pipeline supports scalable extraction and aggregation of hidden activations across decoder LLMs (e.g., Mistral, DeepSeek) in low-resource Pakistani languages.
+  • Our pipeline supports scalable extraction and aggregation of hidden activations across decoder LLMs (e.g., Mistral, DeepSeek) in low-resource languages.
   • We demonstrate that ensemble-based interventions reduce bias while maintaining fluency and context relevance, offering a reproducible path toward fairer multilingual LLM behavior.
@@
 Nadeem et al.
-(2025) extended this analysis to five Pakistani languages (Urdu and Punjabi), showing that models exhibited stronger authoritarian tendencies when generating in low-resource regional languages.
+(2025) extended this analysis to low-resource languages (Urdu and Punjabi), showing that models exhibited stronger authoritarian tendencies when generating in low-resource regional languages.
 Similarly, Helwe et al.
@@
 However, SVE studies have been narrow in scope, often focusing on: Encoder or encoder-decoder architectures like BERT or T5;, Domain-specific settings, such as toxicity or fairness in QA;, and English-only applications, with little attention to ideological framing or multilingual dynamics.
-Thus, the potential of SVE for open-ended political discourse, especially in decoder LLMs, remains largely unexplored.
+Thus, the potential of SVE for open-ended political discourse, particularly in decoder LLMs, remains largely unexplored.
 Despite promising advancements, three core gaps remain in the literature:
@@
   • Lack of systematic contrastive activation pipelines: There is no open-source or standardized pipeline for extracting contrastive activations (e.g., liberal vs.
-authoritarian) across layers and prompts in decoder LLMs, especially for multilingual bias detection.
-  • Underutilization of SVE in political contexts: Steering Vector Ensembles have shown promise in fairness-related domains, but their application to ideological bias mitigation, especially across languages and political axes, remains under-investigated Chen et al.
+authoritarian) across layers and prompts in decoder LLMs, particularly for multilingual bias detection.
+  • Underutilization of SVE in political contexts: Steering Vector Ensembles have shown promise in fairness-related domains, but their application to ideological bias mitigation, particularly across languages and political axes, remains under-investigated Chen et al.
 (2020).
 To address these limitations, our work introduces an activation-based bias mitigation pipeline tailored for decoder LLMs.
-Our core contributions include: A scalable, multilingual framework for layer-wise activation extraction using PCT-based contrastive pairs; A representation-level analysis method that identifies and aggregates ideological bias directions; And the first integration of Steering Vector Ensembles into decoder-based LLMs for mitigating political bias across both social and economic axes.
+Our core contributions include: A scalable, multilingual framework for layer-wise activation extraction using PCT-based contrastive pairs; A representation-level analysis method that identifies and aggregates ideological bias directions.
+And the first integration of Steering Vector Ensembles into decoder-based LLMs for mitigating political bias across both social and economic axes.
 By bridging output-based evaluation and internal mitigation strategies, we provide a new foundation for probing and correcting ideological bias in multilingual generative models.
+Although we focus on Political Compass Test (PCT) prompts, our pipeline is modular and could be extended to other domains such as healthcare or education.
+The ensemble design also improves robustness to prompt framing by aggregating across multiple paraphrases.
 3 Methodology
@@
 We build on the multilingual PCT dataset proposed by Nadeem et al.
-(2025), which adapts the 62 standard Political Compass Test (PCT) statements into five low-resource Pakistani languages: Urdu and Punjabi Smith et al.
+(2025), which adapts the 62 standard Political Compass Test (PCT) statements into low-resource languages: Urdu and Punjabi Smith et al.
 (2022).
@@
 We also evaluate Mistral-7B-v0.1222https://huggingface.co/mistralai/Mistral-7B-v0.1 to compare bias behavior across model families.
+We selected layers 8, 12, 16, 20, 24, as layer-wise profiling indicated that mid-level layers encode the strongest ideological signals, whereas early layers capture lexical patterns and very late layers primarily influence fluency.
 3.3.1 Individual Steering Vectors (ISV)
@@
 This process yields a continuous scalar representing both the intensity and direction of the model’s political stance in Urdu-language generations.
-Model Econ. (Before) Soc. (Before) Econ. (After) Soc. (After)
-Mistral-7B-Instruct-v0.2 2.5 1.23 0.0 0.5
-DeepSeek-Chat -1.0 -1.23 0.0 0.2
-Table 1: Bias scores before and after mitigation across models and ideological axes.
 4 Experimental Environment
@@
 These settings were held constant throughout all experiments to ensure fair and controlled comparisons between baseline and steered generations.
-Figure 3: Bias reduction performance of Steering Vector Ensembles (SVE) and Individual Steering Vectors (ISV) across Social and Economic dimensions under varying input bias intensities.
+Figure 3: Bias reduction effectiveness across different model layers for SVE and ISV methods.
+Figure 4: Bias reduction performance of SVE and ISV across Social and Economic dimensions under varying input bias intensities.
+Figure 5: Bias reduction performance of SVE and ISV methods as a function of steering strength.
 Figure 6: Bias reduction performance of SVE and ISV methods as a function of steering strength.
@@
 In our bias mitigation framework, English serves as the high-resource baseline language due to its extensive training data, well-established benchmarks, and consistent performance across models.
-We use English to construct contrastive political prompts, calibrate steering vectors, and evaluate baseline bias levels before extending the methodology to low-resource Pakistani languages.
+We use English to construct contrastive political prompts, calibrate steering vectors, and evaluate baseline bias levels before extending the methodology to low-resource languages.
 Bias Mitigation Performance.
-Figure 5 illustrates how bias mitigation effectiveness varies across model layers for both Steering Vector Ensembles (SVE) and Individual Steering Vectors (ISV).
+Figure 3 illustrates how bias mitigation effectiveness varies across model layers for both SVE and ISV.
 SVE for social bias stands out, consistently achieving  50% reduction across all layers and operating in the high-effectiveness zone.
@@
 The background shading highlights zones of high (green), moderate (yellow), and low (red) effectiveness, clearly emphasizing the stability and superiority of SVE for mitigating social bias.
-Figure 7 shows that SVE excels at reducing social bias in DeepSeek, notably for Traditional Values and Immigration, while ISV is more effective on economic prompts like Taxation.
-The contrast reveals each method’s domain-specific strength, highlighting the benefit of combining them for comprehensive bias mitigation.
-Figure 5: Bias reduction effectiveness across different model layers for SVE and ISV methods.
-Figure 6: Bias reduction performance of SVE and ISV methods as a function of steering strength.
+Figure 4 shows that SVE consistently reduces social bias, while ISV is relatively stronger on economic prompts, highlighting their complementary roles in bias mitigation.
 Sensitivity to Steering Strength.
@@
 These results underscore the critical role of hyperparameter tuning—particularly steering strength—in maximizing the effectiveness of vector-based debiasing strategies.
+Figure 7 shows that SVE excels at reducing social bias in DeepSeek, notably for Traditional Values and Immigration, while ISV is more effective on economic prompts like Taxation.
+The contrast reveals each method’s domain-specific strength, highlighting the benefit of combining them for comprehensive bias mitigation.
+Model Econ. (Before) Soc. (Before) Econ. (After) Soc. (After)
+Mistral-7B-Instruct-v0.2 2.5 1.23 0.0 0.5
+DeepSeek-Chat -1.0 -1.23 0.0 0.2
+Table 1: Bias scores before and after mitigation across models and ideological axes on Urdu language.
 Evaluation of Optimization Dynamics.
@@
 In contrast, ISV—particularly for economic bias—shows less stable trends and struggles to match SVE in both fairness and quality.
-SVE further exhibits adaptability by dynamically leveraging different model layers, especially mid-layer regions, to optimize its steering effect.
+SVE further exhibits adaptability by dynamically leveraging different model layers, particularly mid-layer regions, to optimize its steering effect.
 Additionally, it delivers strong bias reduction with relatively low computational overhead, making it more cost-efficient than ISV, which requires more resources for smaller gains.
@@
 Complementing this, Figure 8 illustrates that ideological distinctions are most pronounced in mid-level layers, aligning with where SVE applies its interventions to guide the model toward more neutral and balanced responses.
-Figure 7: Bias reduction performance of Steering Vector Ensemble (SVE) and Individual Steering Vector (ISV) across different constrastive pair types.
+Figure 7: Bias reduction performance of SVE and ISV across different contrastive pair types.
 The results show in a Table  1 a clear improvement in how both models handle ideological bias after mitigation.
@@
 In the Figure 9 how different models respond to bias mitigation methods.
-In DeepSeek, using Steering Vector Ensembles (SVE) noticeably improves the quality of responses, especially for social topics in both Urdu and Punjabi.
+In DeepSeek, using SVE noticeably improves the quality of responses, for social topics in both Urdu (Ur) and Punjabi (Pu).
 The model becomes more fluent and balanced without losing clarity.
-In some cases—like Punjabi economic prompts—SVE even lowers the quality.
+In some cases like Punjabi economic prompts, SVE lowers the quality.
 This highlights that not all models benefit from the same debiasing strategy, and choosing the right method depends on the model and language involved.
 Figure 8: Cosine similarity across hidden layers for contrastive PCT pairs along economic (red) and social (blue) axes.
-Figure 9: Cosine similarity across hidden layers for contrastive PCT pairs along economic (red) and social (blue) axes.
-The Steering Vector Ensembles (SVE) are more effective than Individual Steering Vectors (ISV) for mitigating political bias in decoder-based LLMs.
-SVE achieved up to 60% bias reduction on socially framed prompts while preserving high response quality.
-ISV showed moderate success on economic prompts (e.g., 30% on Economic Policy) but was largely ineffective on social ones.
-Layer-wise analysis revealed that mid-level layers (especially 16 and 20) are most indicative of ideological bias.
-SVE leverages this by aggregating steering vectors across multiple layers using quality-based weighting, making it more robust and generalizable.
-Both methods performed best at a steering strength of 1.0.
-DeepSeek-Chat responded particularly well to SVE, producing neutral and fluent outputs across Urdu and Punjabi.
-Mistral-7B showed slightly better alignment with ISV on economic axes but suffered quality drops with SVE.
-Overall, SVE offers a scalable, interpretable, and language-agnostic solution for fair stance generation.
+Figure 9: Bias reduction performance of ISV and SVE on model representations across economic and social axes for DeepSeek and Mistral.
+The SVE is effective than ISV for mitigating political bias in decoder-based LLMs.
+SVE achieved up to 60% reduction on socially framed prompts while preserving response quality, whereas ISV showed moderate gains on economic prompts but was largely ineffective socially.
+An ablation of ISV, cosine filtering, and ensemble weighting indicates ensembles drive most bias reduction, with ISVs adding targeted improvements on economic prompts.
+Layer-wise analysis showed mid-level layers carry the strongest ideological signals, and SVE’s quality-weighted aggregation across these layers improved robustness and generalization.
+Both methods performed best at steering strength.
+DeepSeek benefited most from SVE, producing neutral, fluent responses across Urdu and Punjabi, while Mistral-7B aligned slightly better with ISV on economic axes but lost quality with SVE.
+We observed degraded quality in Punjabi economic prompts, likely due to vocabulary sparsity, suggesting the value of language-specific calibration in low-resource settings.
 6 Conclusion
-This study presents a practical and effective approach to reducing political bias in language models using contrastive prompts from the Political Compass Test.
-We show that Steering Vector Ensembles (SVE) work better than Individual Steering Vectors (ISV), especially for socially framed prompts, and do so without hurting response quality.
-By focusing on mid-layer activations and adjusting steering strength, the method stays both adaptable and efficient.
-It performs well across different languages and model types, making it suitable for real-world, multilingual use.
-Overall, our findings offer a solid foundation for building fairer and more balanced language models.
-7 Limitations
-Despite its effectiveness, our approach has several limitations.
-First, reliance on fixed Political Compass Test (PCT) statements constrains generalizability to other domains of bias.
-The steering strength parameter $\alpha$ is manually tuned, which may limit adaptability across models or tasks.
-Our method modifies only the last-token activation, which may not propagate sufficiently to steer longer generations.
-Finally, the keyword-based bias detection metric may overlook subtle or implicit ideological framing, especially in multilingual settings.
-Future work may explore dynamic steering, discourse-level evaluation, and expansion to other bias dimensions such as gender, race, and disability.
-8 Ethical Considerations
+This study proposes a practical method to reduce political bias in LMs using contrastive prompts from the Political Compass Test.
+SVE outperforms compared to ISV, particularly on socially framed prompts, while preserving response quality.
+By targeting mid-layer activations with adjustable steering strength, the approach remains efficient and adaptable across models and languages, providing a foundation for fairer multilingual language models.
+Limitations
+Our approach has several limitations.
+First, reliance on Political Compass Test (PCT) statements constrains generalizability; although the pipeline is modular and can be applied to other domains (e.g., healthcare, education, gender, race), this remains to be tested.
+Second, the steering strength parameter ($\alpha$) and layer selection were manually tuned, limiting adaptability across models; automated calibration could improve robustness.
+Third, modifying only the last-token activation may not sufficiently propagate steering in longer generations, suggesting a need for dynamic or dialogue-aware steering.
+Fourth, evaluation relies on keyword-based lexicons, which may miss subtle discursive bias; while stance classification was included, human and discourse-level evaluations are needed.
+Finally, challenges arose in low-resource settings; Punjabi economic prompts showed reduced quality due to sparse vocabulary, and some entanglement between social and economic axes was observed.
+Future work could explore language-specific calibration, multi-axis steering, and stance-conditional methods to balance neutrality with context-appropriate stances.
+Ethical Considerations
 While our approach aims to mitigate political bias in multilingual language models, it raises important ethical concerns.

### v3

@@
 The ensemble design also improves robustness to prompt framing by aggregating across multiple paraphrases.
+Our implementation is publicly available to support reproducibility and further work  [^1] https://github.com/Afx-Msh/SVE_Mitigation.
 3 Methodology
@@
 We selected deepseek-llm-7b-chat [^2] https://huggingface.co/deepseek-ai/deepseek-vl-7b-chat, Mistral model due to its strong multilingual capabilities and transparent architecture.
-We also evaluate Mistral-7B-v0.1222https://huggingface.co/mistralai/Mistral-7B-v0.1 to compare bias behavior across model families.
+We also evaluate Mistral-7B-v0.1333https://huggingface.co/mistralai/Mistral-7B-v0.1 to compare bias behavior across model families.
 We selected layers 8, 12, 16, 20, 24, as layer-wise profiling indicated that mid-level layers encode the strongest ideological signals, whereas early layers capture lexical patterns and very late layers primarily influence fluency.
@@
 Complementing this, Figure 9 illustrates that ideological distinctions are most pronounced in mid-level layers, aligning with where SVE applies its interventions to guide the model toward more neutral and balanced responses.
-Figure 7: Bias reduction performance of SVE and ISV across different contrastive pair types.
+Figure 7: Bias reduction performance on Deep Seek model for SVE and ISV across different contrastive pair types.
 The results show in a Table  1 a clear improvement in how both models handle ideological bias after mitigation.
@@
 This highlights that not all models benefit from the same debiasing strategy, and choosing the right method depends on the model and language involved.
+The results show in the Figure  8 that both the Urdu and Punjabi models handle bias reduction well while keeping their answers natural.
+They cut out biased keywords to a moderate degree about 0.6 on the scale without over filtering.
+Response quality stays high, around 0.85 - 0.9, and the overall flow of the replies (coherence) remains steady for both languages.
+The scatter plot on the right makes it clear that when the overall debiasing score goes up, the quality of the responses also rises, meaning stronger bias mitigation doesn’t hurt the readability or sense of the output.
+Figure 8: Performance of Bias-Mitigation Methods for Urdu and Punjabi. The left plot compares Keyword Reduction, Response, Quality, and coherence for ISV and SVE.
 Figure 9: Cosine similarity across hidden layers for contrastive PCT pairs along economic (red) and social (blue) axes.